OpenStudy (anonymous):

**WILL MEDAL** Which ordered pair is the solution to the system of equations? 2x+y=-8 x-y=-4 A. (-3, -2) B. (-1, 3) C. (1, 5) D. (-4, 0)

OpenStudy (anonymous):

c

OpenStudy (anonymous):

Thx

OpenStudy (anonymous):

2x + y = -8 + x - y = -4 ------------ 3x = -12 x = -4 2(-4) + y = -8 -8 + y = -8 y = 0 D) (-4, 0)

OpenStudy (haleyelizabeth2017):

It's actually d....

OpenStudy (anonymous):

Oh ok thx guys